OpenStudy (anonymous):

The lodge at the Three Lakes Resort is offering the following two options for a holiday weekend. Option 1: 3 nights including six meals for $132.00 Option 2: 3 nights including two meals for $109.00 At these rates, what is the cost of one night's lodging?

OpenStudy (anonymous):

Change each sentence into an algebraic equation. 3N+6M=132 3N+2M=109

OpenStudy (anonymous):

Of course, you have to assume the nightly rate is the same and the per meal cost is the same.

OpenStudy (anonymous):

Using the first equation: 3N=132-6M N=(132-6M)/3 Substitute into the second equation, the value of N 3(132-6M)/3 + 2M = 109 132-6M+2M=109 4M=23 M=5.75 Substitute the value M into either equation: 3N+2(5.75)=109 3N=109-11.50 3N=97.50 N=32.50

OpenStudy (anonymous):

or we could reason as follows: since the nights are the same you know that 4 meals cost $132-$109=$23 therefore two meals cost $11.5 (rather cheap) and so 3 nights cost $109-$11.5=$97.50 and so one night cost $32.50